# On the Convergence Rate of LoRA Gradient Descent

### v1

@@ 3.2 Results
Theorem 3 states that LoRA gradient descent converges to a stationary point at $O(\frac{1}{\log T})$ rate. If $\lVert V_{t}\rVert$ is uniformly bounded, the rate improves to $O(1/T)$. This matches the standard convergence rate of gradient descent on Lipschitz smooth functions and the rates derived by existing works that analyze LoRA under the bounded parameter assumption, which essentially forces $\mathcal{J}$ to be Lipschitz smooth. Our results demonstrate that the behavior of LoRA is governed by complex interactions between $\eta_{t}$, $\lVert V_{t}\rVert$, and $\lVert\nabla\mathcal{J}(V_{t})\rVert$, which slows convergence. However, even if $\lVert V_{t}\rVert$ grows and forces $\eta_{t}$ to decrease at each iteration, the algorithm still converges; a simple explanation is that if $\lVert V_{t}\rVert$ grows, $\eta_{t}$ decreases, which slows the growth of $\lVert V_{t}\rVert$.
4 Discussion
@@
Our results yield several interesting insights into the behavior of LoRA. First, the choice of extractor matrices $E_{1}$, $E_{2}$, has little to no impact on the proof, suggesting that other subsets of the $VV^{T}$ matrix could be extracted as alternate parameter-efficient reparametrizations to yield similar results. Moreover, the $VV^{T}$ form is exactly the symmetric Burer-Monteiro parametrization [burer_nonlinear_2003], and the results of our analysis can be trivially extended to show that gradient descent with Burer-Monteiro also converges to stationary points for general Lipschitz smooth functions. Second, the convergence rate does not explicitly depend on the chosen LoRA rank $r$, except that it is $O(\frac{1}{\lVert V_{0}\rVert^{2}})$. This is expected because $r$ just controls the dimension of the gradient descent algorithm, which is a dimension-free algorithm. Finally, our results shed light onto the complex relationship between the parameter norm and the LoRA training dynamics, which has been alluded to in past work. In particular, we observe that the convergence of LoRA is slowed if the iterates are progressing away from the origin, and it is accelerated if they are moving towards the origin. This “position dependency” is not typical of gradient algorithms, and may be due to the fact that the LoRA reparametrization creates a stationary point at $V=0$ regardless of the structure of the original function. Consequently, LoRA may converge to the origin even if the original global minima is arbitrarily far away, explaining the divergence in behavior between LoRA and full-rank fine-tuning.
@@ A.1 Helper Lemmas
For the proof of Lemma 4, see Section 2.5 of [hardy_inequalities_1934].
Lemma 5.
@@ A.3 Proof of Lemma 2
We apply the descent lemma (Lemma 1) with $V_{1}=V_{t}$, $V_{2}=V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})$, which yields
@@
Therefore, from (14) of Lemma 4, we have $\frac{1}{a+b}\leq\frac{1}{2\sqrt{ab}}$ for $a,b>0$, such that $\eta_{t}$ satisfies the following bound,
@@
For the next bound, by (13) of Lemma 4 with $w_{1}=1$, $w_{2}=2$, $x_{1}=5\sqrt{2}\lVert V_{t}\rVert^{2}$, $x_{2}=\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$, we have
@@
In conclusion, with the choice of $\eta_{t}$ in (8), we can substitute the derived bounds on $\eta_{t}$ into Lemma 1, achieving descent in one step.
@@ A.4 Proof of Theorem 3
From Lemma 2, we know that if
@@
So to show convergence to a stationary point, we need to show that the sum of learning rates $\sum_{t=0}^{\infty}\eta_{t}$ diverges. From Lipschitz smoothness of $\mathcal{L}$, we have for all $W\in\mathbb{R}^{m\times n}$ (See Lemma 2.28 in [garrigos2024handbookconvergencetheoremsstochastic]),
$\lVert\nabla\mathcal{L}(W)\rVert^{2}\leq$ $2L(\mathcal{L}(W)-\mathcal{L}^{*})$
$\lVert\nabla\mathcal{L}(W)\rVert\leq$ $(2L(\mathcal{L}(W)-\mathcal{L}^{*}))^{1/2}$
@@
To show the divergence of $\sum_{t=0}^{\infty}\eta_{t}$, it is sufficient to show the divergence of the series $\sum_{t=0}^{\infty}\min\{\frac{1}{D_{t}},1\}$. Moreover, by the contrapositive of Lemma 6, we only need to show divergence of $\sum_{t=0}^{\infty}\frac{1}{D_{t}}$. We first show that we achieve divergence if the adapter norms are bounded. If $\lVert V_{t}\rVert\leq C$ for all $t$, then $\eta_{t}$ is lower bounded by a nonzero constant $\eta>0$
@@
and we prove (12) of Theorem 3 as follows

### v2

@@ 3.2 Results
Theorem 3.5 states that LoRA gradient descent converges to a stationary point at $O(\frac{1}{\log T})$ rate. If $\lVert V_{t}\rVert$ is uniformly bounded, the rate improves to $O(1/T)$. This matches the standard convergence rate of gradient descent on Lipschitz smooth functions and the rates derived by existing works that analyze LoRA under the bounded parameter assumption, which essentially forces $\mathcal{J}$ to be Lipschitz smooth. Our results demonstrate that the behavior of LoRA is governed by complex interactions between $\eta_{t}$, $\lVert V_{t}\rVert$, and $\lVert\nabla\mathcal{J}(V_{t})\rVert$, which slows convergence. However, even if $\lVert V_{t}\rVert$ grows and forces $\eta_{t}$ to decrease at each iteration, the algorithm still converges; a simple explanation is that if $\lVert V_{t}\rVert$ grows, $\eta_{t}$ decreases, which slows the growth of $\lVert V_{t}\rVert$.
Extension to multiple matrices.
The above theory assumes the loss functions is parametrized by a single weight matrix $W$. In practice, deep neural networks are parametrized by multiple weight matrices of potentially different sizes, each with its own LoRA adapter matrices. Our analysis naturally extends to this setting, which we illustrate (without loss of generality) with the following example of two matrices. Let $\mathcal{L}(W_{1},W_{2})\in\mathbb{R}$ denote the loss as a function of $W_{1}\in\mathbb{R}^{m_{1}\times n_{1}}$ and $W_{2}\in\mathbb{R}^{m_{2}\times n_{2}}$. For any matrix $M\in\mathbb{R}^{(m_{1}+m_{2})\times(n_{1}+n_{2})}$, define the linear operators $E_{11}(M)\in\mathbb{R}^{m_{1}\times n_{1}}$ and $E_{22}(M)\in\mathbb{R}^{m_{2}\times n_{2}}$ as follows,
$E_{11}(M)=\begin{bmatrix}I_{m_{1}}&0_{m_{1}\times m_{2}}\end{bmatrix}M\begin{bmatrix}I_{n_{1}}\\ 0_{n_{2}\times n_{1}}\end{bmatrix}$,
$E_{22}(M)=\begin{bmatrix}0_{m_{2}\times m_{1}}\times I_{m_{2}}\end{bmatrix}M\begin{bmatrix}0_{n_{1}\times n_{2}}\\ I_{n_{2}}\end{bmatrix}$,
where $E_{11}$ extracts the top left $m_{1}\times n_{1}$ block and $E_{22}$ extracts the bottom right $m_{2}\times n_{2}$ block of $M$. We can define a new function $\tilde{\mathcal{L}}:\mathbb{R}^{(m_{1}+m_{2})\times(n_{1}+n_{2})}\to\mathbb{R}$ such that
$\tilde{\mathcal{L}}(M)=\mathcal{L}(E_{11}(M),E_{22}(M)),$
and the gradient $\nabla\tilde{\mathcal{L}}(M)$ is equal to the following matrix,
$\begin{bmatrix}\nabla_{W_{1}}\mathcal{L}(E_{11}(M),E_{22}(M))&0\\ 0&\nabla_{W_{2}}\mathcal{L}(E_{11}(M),E_{22}(M))\end{bmatrix}$.
This follows from the fact that by construction, the terms in the top right and bottom left blocks of $M$ have no impact on the value of $\tilde{\mathcal{L}}$, and their derivatives are zero.
Overloading notation, we define the $\ell^{2}$ product norm on $\mathbb{R}^{m_{1}\times n_{1}}\times\mathbb{R}^{m_{2}\times n_{2}}$ as follows, for matrices $W_{1}\in\mathbb{R}^{m_{1}\times n_{1}}$ and $W_{2}\in\mathbb{R}^{m_{2}\times n_{2}}$,
$\lVert(W_{1},W_{2})\rVert^{2}_{F}=\lVert W_{1}\rVert^{2}_{F}+\lVert W_{2}\rVert^{2}_{F}.$
Then we can show that if $\mathcal{L}$ is Lipschitz smooth, $\tilde{\mathcal{L}}$ is also Lipschitz smooth.
Lemma 3.6.
Suppose $\mathcal{L}$ is $L$-Lipschitz smooth such that for all $W_{1},W_{1}^{\prime}\in\mathbb{R}^{m_{1}\times n_{1}}$, and $W_{2},W_{2}^{\prime}\in\mathbb{R}^{m_{2}\times n_{2}}$,
$\lVert\nabla\mathcal{L}(W_{1},W_{2})-\nabla\mathcal{L}(W_{1}^{\prime},W_{2}^{\prime})\rVert\leq L\lVert(W_{1}-W_{1}^{\prime},W_{2}-W_{2}^{\prime})\rVert.$
Then $\tilde{\mathcal{L}}$ is $2L$-Lipschitz smooth.
Proof. See Appendix A.5.
Finally, we reparametrize $W_{1}=B_{1}A_{1}$ and $W_{2}=B_{2}A_{2}$, where $B_{1}\in\mathbb{R}^{m_{1}\times r}$, $A_{1}\in\mathbb{R}^{r\times n_{1}}$, $B_{2}\in\mathbb{R}^{m_{2}\times r}$, $A_{2}\in\mathbb{R}^{r\times n_{1}}$ represent the LoRA adapter matrices. We can construct the larger low-rank matrices $B=[B_{1}^{T},B_{2}^{T}]^{T}$ and $A=[A_{1},A_{2}]$, such that
$M=BA=\begin{bmatrix}B_{1}A_{1}&B_{1}A_{2}\\ B_{2}A_{1}&B_{2}A_{2}\end{bmatrix}$.
Then $B_{1}A_{1}=E_{11}(BA)$ and $B_{2}A_{2}=E_{22}(BA)$, and gradient descent on $\tilde{L}(BA)$ with respect to $B$ and $A$ is equivalent to LoRA gradient descent on $\mathcal{L}$. Since $\tilde{L}$ is Lipschitz smooth, we can directly apply the existing analysis to this setting.
Figure 1: Training logistic regression model on embeddings of the CIFAR-10 dataset with constant, adaptive ($\eta^{adapt}$ and $\eta^{adapt2}$), and normalized ($\eta^{norm}$) learning rates. For clarity, the moving averages (window size $10$ and $200$ for a. and b. respectively) of the batch loss are plotted.
3.3 Discussion
Our theoretical results yield several interesting insights into the behavior of LoRA. First, the choice of extractor matrices $E_{1}$, $E_{2}$, has little to no impact on the proof, suggesting that other subsets of the $VV^{T}$ matrix could be extracted as alternate parameter-efficient reparametrizations to yield similar results. Moreover, the $VV^{T}$ form is exactly the symmetric Burer-Monteiro parametrization (Burer and Monteiro, 2003), and the results of our analysis can be trivially extended to show that gradient descent with Burer-Monteiro also converges to stationary points for general Lipschitz smooth functions. Second, the convergence rate does not explicitly depend on the chosen LoRA rank $r$, except that it is $O(\frac{1}{\lVert V_{0}\rVert^{2}})$. This is expected because $r$ just controls the dimension of the problem, but gradient descent is a dimension-free algorithm. Finally, due to the $\lVert V_{t}\rVert^{2}$ term in the denominator of $\eta_{t}$, the convergence of LoRA is slowed if the iterates are progressing away from the origin, and it is accelerated if they are moving towards the origin. This “position dependency” is not typical of gradient algorithms, and may be due to the fact that the LoRA reparametrization creates a stationary point at $V=0$ regardless of the structure of the original function. Consequently, LoRA may converge to the origin even if the original global minima is arbitrarily far away, explaining the observed divergence in behavior between LoRA and full-rank fine-tuning (Shuttleworth et al., 2025; Biderman et al., 2024).
Figure 2: Training ResNet-18 model on the CIFAR-10 dataset with constant, adaptive ($\eta^{adapt2}$) and normalized ($\eta^{norm}$) learning rates. For clarity, the moving averages (window size $50$) of the batch loss and gradient norm are plotted.
4 Experiments
We conduct experiments to investigate how the choice of learning rate and LoRA reparametrization affects training convergence in practice. We perform image classification on the CIFAR-10 dataset (Krizhevsky, 2009) with 10 classes, using the cross-entropy loss function. We first consider logistic regression, since the loss function is known to be Lipschitz smooth, and we train the model on embeddings of the CIFAR-10 images, produced by a ResNet-18 model (He et al., 2016) pretrained on ImageNet-1k (Deng et al., 2009). We then train a ResNet-18 model directly on the CIFAR-10 dataset, with LoRA implemented on the convolutional layers, and batch normalization layers disabled because they can have complex interactions with the learning rate. For each experimental setting, we replace the model weight matrices with low-rank approximations ($r=4$ and $r=20$) and we train with mini-batch SGD with a large batch size ($b=512$) to better capture our analysis of gradient descent. The model weights are first initialized randomly and frozen, and then the adapter matrices are initialized in the standard approach with $A$ as a random matrix, and $B$ as a zero matrix. For additional experimental details, including the anonymous GitHub repository, see Appendix B.
We test three types of learning rate schemes: constant, adaptive, and normalized. We first consider an adaptive learning rate $\eta^{adapt}$ defined as follows,
$\eta^{adapt}_{t}=\frac{\alpha}{\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert}.$
[…]
$\eta_{t}^{adapt2}=\frac{\alpha}{\lVert V_{t}\rVert^{2}+\sqrt{\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})}}.$
Finally, we also consider a normalized learning rate $\eta_{norm}$, defined as follows,
$\eta^{norm}_{t}=\frac{\alpha}{\lVert\nabla\mathcal{L}(V_{t})\rVert^{1/2}}.$
This inverse relationship between $\eta_{t}$ and $\lVert\nabla\mathcal{L}(V_{t})\rVert^{1/2}$ is required for minimizing higher order terms in the descent lemma, as stated in (21) and (23) in Appendix A.3. In practice, we compute $\eta^{adapt}_{t}$, $\eta^{adapt2}_{t}$ and $\eta^{norm}_{t}$ using the batch gradient and loss values at training step $t$.
Figure 1 displays the results of the logistic regression experiments. Figures 1a and 1b demonstrate that training with the nonconstant learning rates $\eta_{t}^{adapt}$, $\eta_{t}^{adapt2}$, or $\eta_{t}^{norm}$ exhibits faster and more stable convergence than training with a constant learning rate in the same range of values. If the constant learning rate is too high, training becomes unstable, but if it is too low, convergence slows. Figure 1c displays the values of $\eta^{adapt}$, $\eta^{adapt2}$ and $\eta^{norm}$ over the training time, demonstrating that $\eta^{adapt}_{t}$ and $\eta^{adapt2}_{t}$ are closely correlated early on, but diverge as the model converges.
Figure 2 displays the results of the ResNet-18 experiments using $\eta_{t}^{adapt2}$ and $\eta_{t}^{norm}$. Both the adaptive and normalized learning rates significantly stabilize training compared to constant learning rate while maintaining fast convergence, with $\eta^{norm}$ outperforming $\eta^{adapt2}$. Therefore, $\eta^{adapt2}$ and $\eta^{norm}$ are practical and computationally efficient methods for improving LoRA convergence on large neural networks.
Figure 3: Training logistic regression model on embeddings of the CIFAR-10 dataset with constant, adaptive ($\eta^{adapt}$ and $\eta^{adapt2}$), and normalized ($\eta^{norm}$) learning rates. For clarity, the moving averages (window size $30$) of the batch gradient are plotted.
Figures 3a and 2c display the progress of $\lVert V_{t}\rVert$ for various learning rates, which in combination with Figures 3b and 2d, suggests that the model is initialized in a flat region close to the origin, where the gradient norm is small. This aligns with the fact that LoRA always creates a stationary point at the origin. Our results suggest that a large initial learning rate is necessary for accelerating the iterates out of this region.
Ultimately, our results demonstrate that $\eta^{adapt}$, $\eta^{adapt2}$, and $\eta^{norm}$ can improve LoRA training by taking advantage of the structural changes introduced by the low-rank reparametrization. These learning rate schemes start at higher values when the gradient norm is low and decrease over time, thereby accelerating training through the plateau around initialization and stabilizing it when the gradient norm is large. In particular, $\eta^{adapt}$ is most aligned with theory, while $\eta^{adapt2}$ and $\eta^{norm}$ are more appropriate for practical large-scale implementations.
5 Conclusion
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Helper Lemmas
For the proof of Lemma A.1, see Section 2.5 of (Hardy et al., 1934).
Lemma A.2.
@@
We apply the descent lemma (Lemma 3.3) with $V_{1}=V_{t}$, $V_{2}=V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})$, which yields
@@
Therefore, from (14) of Lemma A.1, we have $\frac{1}{a+b}\leq\frac{1}{2\sqrt{ab}}$ for $a,b>0$, such that $\eta_{t}$ satisfies the following bound,
@@
For the next bound, by (13) of Lemma A.1 with $w_{1}=1$, $w_{2}=2$, $x_{1}=5\sqrt{2}\lVert V_{t}\rVert^{2}$, $x_{2}=\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$, we have
@@
In conclusion, with the choice of $\eta_{t}$ in (8), we can substitute the derived bounds on $\eta_{t}$ into Lemma 3.3, achieving descent in one step.
@@
From Lemma 3.4, we know that if
@@
So to show convergence to a stationary point, we need to show that the sum of learning rates $\sum_{t=0}^{\infty}\eta_{t}$ diverges. From Lipschitz smoothness of $\mathcal{L}$, we have for all $W\in\mathbb{R}^{m\times n}$ (See Lemma 2.28 in (Garrigos and Gower, 2024)),
$\lVert\nabla\mathcal{L}(W)\rVert^{2}\leq 2L(\mathcal{L}(W)-\mathcal{L}^{*}),$ (24)
which allows us to bound the gradient as follows,
$\lVert\nabla\mathcal{L}(W)\rVert\leq$ $(2L(\mathcal{L}(W)-\mathcal{L}^{*}))^{1/2},$
@@
To show the divergence of $\sum_{t=0}^{\infty}\eta_{t}$, it is sufficient to show the divergence of the series $\sum_{t=0}^{\infty}\min\{\frac{1}{D_{t}},1\}$. Moreover, by the contrapositive of Lemma A.3, we only need to show divergence of $\sum_{t=0}^{\infty}\frac{1}{D_{t}}$. We first show that we achieve divergence if the adapter norms are bounded. If $\lVert V_{t}\rVert\leq C$ for all $t$, then $\eta_{t}$ is lower bounded by a nonzero constant $\eta>0$
@@
and we prove (12) of Theorem 3.5 as follows
@@
A.5 Proof of Lemma 3.6
Suppose $M,M^{\prime}\in\mathbb{R}^{(m_{1}+m_{2})\times(n_{1}+n_{2})}$. Then we have
$\lVert\nabla\tilde{L}(M)-\nabla\tilde{L}(M^{\prime})\rVert^{2}=$ $\lVert\nabla_{W_{1}}\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla_{W_{1}}\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert^{2},$
$+\lVert\nabla_{W_{2}}\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla_{W_{2}}\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert^{2},$
$=$ $\lVert\nabla\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert^{2},$
$\lVert\nabla\tilde{L}(M)-\nabla\tilde{L}(M^{\prime})\rVert=$ $\lVert\nabla\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert,$
$\leq$ $L\lVert(E_{11}(M),E_{22}(M))-(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert,$
$=$ $L\lVert(E_{11}(M-M^{\prime}),E_{22}(M-M^{\prime}))\rVert,$
$\leq$ $L\lVert E_{11}(M-M^{\prime})\rVert+L\lVert E_{22}(M-M^{\prime})\rVert\leq 2L\lVert M-M^{\prime}\rVert,$
where the last step follows from similar logic as Lemma A.2.
Table 1: Experiment hyperparameters.
Hyperparameter Logistic regression with ResNet-18 embeddings ResNet-18
Batch size 512 512
LoRA rank 4 20
Data input dimension 512 $32\times 32\times 3$
Epochs 60 250
$\alpha^{adapt}$ 1
$\alpha^{adapt2}$ 0.8 15
$\alpha^{norm}$ 0.06 0.1
Constant learning rate range [0.03, 0.05, 0.07] [0.03, 0.04, 0.045]
Appendix B Experimental Details
In the following, we provide additional details about our experimental setup. As mentioned in the paper, for the ResNet-18 experiments we turn off BatchNorm layers because they effectively change the learning rate. This is achieved by setting the model to evaluation mode with model.eval() before training. Table 1 displays the hyperparameters used in our experimental results. Code is open-sourced at the anonymous Github repository https://anonymous.4open.science/r/lora-C709/.
Learning rate.
While standard LoRA implementations tend to scale by both a scaling factor and $\frac{1}{r}$, for simplicity we do not scale by the rank and just set the learning rate as a whole. To determine the range of learning rates to test in our experiments, we use a cyclical learning rate finder as introduced in (Smith, 2017) and further described in (Ahmed, 2025). We start with a small learning rate and train the model for a single epoch, repeating this with exponentially increasing learning rates. We plot the loss over the learning rates and pick learning rates in the range where the loss decreases the fastest. The resulting plots are available in Figure 4.
Figure 4: Learning rate selection using a cyclical finder for both experimental settings.
Hardware and software.
All experiments were run using PyTorch 2.5.0 and CUDA 12.1, on an Intel(R) Xeon(R) Silver 4208 CPU (2.10GHz) with an NVIDIA RTX A6000 GPU.